# The shocks in Josephson transmission line revisited

### 27 Dec 2023 v3

@@ Abstract
We continue our previous studies of the shocks in the lossy Josephson transmission line (JTL). The paper consists of two parts. In the first part we analyse the scattering of the ”sound’ (small amplitude small wave vector harmonic wave) on the shock wave and calculate the reflection and the transmission coefficients. In the second part we show that the kinks, which we previously studied only in the lossless JTL, exist also in the lossy JTL and study the similarities and the dissimilarities between the shocks and the kinks there. We find the cases, when the nonlinear equation, describing the weak shocks and kinks can be integrated in terms of elementary functions. We also show that the profile of the shock in the lossy JTL demonstrates soliton-like features if the losses are weak.
@@ I Introduction
We studied previously the shock waves in the lossy Josephson transmission line (JTL) JTL kogan1 ; kogan2 ; kogan3 and kinks (and solitons) in the lossless (actually, without any shunting at al) JTL kogan2 ; kogan3 . The present work had several aims. First we would like to analyse the interaction between the ”sound” (small amplitude small wave vector harmonic wave) and the shock wave. Second we would like to establish the relation between the shock waves and the kinks. And third, we would like to additionally study the weak waves, and, in particular, to look for the cases when the nonlinear equation, describing such waves in the JTL, can be integrated in terms of elementary functions.
@@ IV.1 The travelling waves
Excluding $V$ from (IV.1) and integrating the resulting equation we obtain closed equation for $\varphi$
$\frac{d^{2}\varphi}{d\tilde{\tau}^{2}}+\gamma\frac{d\varphi}{d\tilde{\tau}}=\overline{U}^{2}\varphi-\sin\varphi+F,$ (26)
where $\tilde{\tau}\equiv\tau\sqrt{C/C_{J}}=t/\sqrt{L_{J}C_{J}}$, $\gamma\equiv\sqrt{L_{J}/C_{J}}\big{/}R_{J}$ is the damping coefficient, and $F$ is the constant of integration. Taking into account the boundary conditions (23) and the equation for the shock speed (9), we can write down (26) as
$\frac{d^{2}\varphi}{d\tilde{\tau}^{2}}+\gamma\frac{d\varphi}{d\tilde{\tau}}+\sin\varphi=\overline{I}(\varphi),$ (27)
where
$\overline{I}(\varphi)\equiv\frac{(\varphi-\varphi_{2})\sin\varphi_{1}-(\varphi-\varphi_{1})\sin\varphi_{2}}{\varphi_{1}-\varphi_{2}},$ (28)
which reminds the equation
@@
Alternatively (26) can be written down as
@@
Because of the invariance of the system when all phases are shifted by $2\pi$, it is enough to consider $\varphi_{1}\in(-\pi,\pi)$. Because of the condition (6), we should consider only $\varphi_{1}\in(-\pi/2,\pi/2)$, and because the physics is obviously symmetric with respect to simultaneous inversion of all phases $\varphi\to-\varphi$, in the following we consider explicitly only $\varphi_{1}\in(0,\pi/2)$.
@@
We considered previously the potential maximum at $\varphi_{2}$ case only in the absence of shunting kogan2 . We called such travelling waves the kinks. Now we understand that similar kinks exist also in the lossy JTL (for $-\varphi_{1}<\varphi_{2}<\varphi_{1}/2$). Looking at Fig. 5 (right), presenting the potential for the kink, we realize, that since the particle stops at the unstable equilibrium point, for kink to exist, the fine tuning of $\gamma$ is necessary. Saying it in different words, for a given $\varphi_{1}$ and given $\gamma$, only the kink with the definite value of $\varphi_{2}$ can exist. In particular, in the absence of losses ($\gamma=0$) only the kinks with $\varphi_{2}=-\varphi_{1}$, are possible kogan2 .
@@
$\frac{\sin\varphi_{1}-\sin\varphi_{2}}{\varphi_{1}-\varphi_{2}}-\cos\varphi_{2}=0.$ (34)
@@
IV.2.1 The shocks speed vs. the kinks speed
@@
From the inequalities (35) follows that a shock can not split into two shocks moving in the same direction. Actually we can make even stronger statement: two shocks moving in the same direction will merge. In fact, let there is the first shock $\varphi_{2}\leftarrow\varphi_{0}$ and the second shock $\varphi_{0}leftarrow\varphi_{1}$ ahead of it. Because of inequalities (35) the speed of the first shock is larger, and the speed of the second shock is smaller that $\overline{u}(\varphi_{0})$. The statement is proved. Note that due to a one-dimensional nature of our problem we don’t have to consider the corrugation instability of the shock wave dyakov1 ; lubchich ; semenko ; robinet ; bates .
@@
IV.3 Elementary solutions for weak shocks
For weak wave, characterized by the condition $\varphi_{1}-\varphi_{2}\ll 1$ (for the sake of definiteness we have chosen $\varphi_{1}>0$), the r.h.s. of (30) can be approximated as
@@ IV.3 Elementary solutions for weak shocks
Note that for the kink, $-\varphi_{3}$ is between $\varphi_{1}$ and $\varphi_{2}$, for the shock, $-\varphi_{3}$ is outside this interval. As the result, (30) can be simplified to the damped Helmholtz-Duffing (dHD) equation
@@
$\varphi(\tilde{\tau})=\varphi_{2}+\frac{\varphi_{1}-\varphi_{2}}{e^{\beta\tilde{\tau}}+1},$ (41)
@@
$\beta=\sqrt{\frac{\alpha}{2}}(\varphi_{1}-\varphi_{2})=\sqrt{\frac{\cos\overline{\varphi}}{12}}(\varphi_{1}-\varphi_{2}).$ (42)
A simple example: for $\varphi_{2}=-\varphi_{1}$, Eqs. (40), (41) and (42) give $\gamma=0$ and
$\varphi(\tilde{\tau})=-\varphi_{1}\tanh\left(\frac{\varphi_{1}\tilde{\tau}}{2\sqrt{3}}\right)$ (43)
(compare with Eq. (39) from Ref. kogan2 ).
If the assumption which lead to Eq. (39) is strengthened to $\varphi_{1}-\varphi_{2}\ll\tan\overline{\varphi}$, there also exists another elementary solution of Eq. (39). It is shown in Appendix D, that under the condition
$\gamma=5\sqrt{\frac{(\varphi_{1}-\varphi_{2})\sin\overline{\varphi}}{12}},$ (44)
the solution of of (39) satisfying the boundary conditions (23) is
$\varphi(\tilde{\tau})=\varphi_{2}+\frac{\varphi_{1}-\varphi_{2}}{\left(e^{\gamma\tilde{\tau}/5}+1\right)^{2}}.$ (45)
Looking at Eq. (39) we realise that for $\varphi_{1}-\varphi_{2}\ll\tan\overline{\varphi}$, the condition (40) corresponds to strong damping and the condition (44) – to critical damping (see Appendix D).
IV.4 The ”solitons” within the shocks
In our previous publication kogan2 we studied the localized travelling waves existing in the JTL in the absence of damping (the kinks and the solitons) and in the presence of damping (the shocks). In Section IV.2 we have shown that the kinks exist also in the presence of the damping and are connected with the shocks. So what about the solitons? Before answering this question let me remind what the solitons in the JTL are kogan2 . Let us look at the Fig. 4. In the absence of the damping ($R_{J}=\infty\rightarrow\gamma=0$) the motion of the fictitious particle starting and ending at $\varphi=\varphi_{1}$ is possible. This corresponds to the boundary conditions (23) with $\varphi_{2}=\varphi_{1}$ (the soliton).
Now let us switch on the damping. We understand that the solitons don’t exist any more, because the particle can’t return to its initial position. However, for weak damping, the particle after leaving the initial equilibrium position $\varphi_{1}$ will nearly return to this position after reflection from the opposite wall of the potential well (see Fig. 4). The part of the shock described by that motion will look very much like the soliton. The point of the reflection $\varphi_{0}$ can be found from the equation
$\Pi(\sin\varphi_{0})\approx\Pi(\sin\varphi_{1}).$ (46)
Note that such ”soliton” will be followed by the other ones, only (because of the continuous decrease of energy of the particle) the distance between the successive ”solitons” will be decreasing and each successive ”soliton” will look less and less like the real soliton.
@@ V Conclusions
We established the relation between the shocks existing in the lossy JTL and the kinks which, as we now understand, exist both in the lossy and in the lossless JTL. We have shown how the existing in the lossless JTL solitons manifest themselves in the profile of the shock, propagating in the JTL with weak losses.
We found the cases when nonlinear equation describing weak shocks (kinks) can be integrated in terms of elementary functions.
@@ Appendix A The JTL composed of superconducting grains
$m\frac{d^{2}y_{n}}{d\tau^{2}}=-\frac{\partial}{\partial y_{n}}\left[U\left(y_{n-1}-y_{n}\right)+U\left(y_{n+1}-y_{n}\right)\right]-\gamma\frac{dy_{n}}{d\tau},$
@@
$\frac{d^{2}\varphi_{n}}{d\tau^{2}}+\gamma\frac{d\varphi_{n}}{d\tau}+\sin\varphi_{n}-\frac{1}{a^{2}}\left(\varphi_{n-1}-2\varphi_{n}+\varphi_{n+1}\right)=I/I_{c},$
@@ Appendix B The continuum and the discrete JTL
Lossless JTL clearly corresponds to $R_{J}=\infty$. In addition, while talking about the lossless system in the present paper, we had in mind Eq. (B) containing only the second term in the l.h.s.. Previously kogan2 , we considered unshunted JTL in the quasi-continuum approximation, that is Eq. (B) with only the first term in the l.h.s.. However, because of the similarity of the terms, the kinks and the solitons obtained in in the framework of these two considerations are qualitatively very similar. (The quantitative differences do exist. Thus the equation for the ”soliton” velocity (47) is different from the equation for the soliton velocity (36) from Ref. kogan2 ).
@@
Appendix D Elementary particular solutions of the generalized dHD equation
@@
$p_{\tau\tau}+\gamma p_{\tau}=\alpha p\left(p^{n}-p_{1}\right)\left(p^{n}+p_{3}\right),$ (61)
where $n$ is an integer, $\gamma$, $\alpha$, $p_{1},p_{3}$ are positive constants ($p=0$ is the stable fixed point and $p_{1},p_{3}$ are unstable fixed points.)
@@
$-\alpha(p_{1}-p_{3})e^{-nm\tau}w^{n+1}-\left(m^{2}-m\gamma+\alpha p_{1}p_{3}\right)w.$ (63)
@@
$n^{2}m^{2}w_{zz}+n\left[(n+1)m^{2}-\alpha p_{1}p_{3}\right]\frac{w_{z}}{z}$
$=\alpha w^{2n+1}-\alpha(p_{1}-p_{3})\frac{w^{n+1}}{z}.$ (66)
Note that (D) is just Eq. (61), only written down in the new variables.

### 28 Feb 2024 v6

@@ Abstract
We continue our previous studies of the shocks in the lossy Josephson transmission line (JTL). The paper consists of two parts. In the first part we analyse the scattering of the ”sound’ (small amplitude small wave vector harmonic wave) on the shock wave and calculate the reflection and the transmission coefficients. In the second part we show that the kinks, which we previously studied only in the lossless JTL, exist also in the lossy JTL and study the similarities and the dissimilarities between the shocks and the kinks there. We find that the nonlinear equation describing the weak kinks and the weak shocks can be integrated (in particular cases) in terms of elementary functions. We also show that the profile of the shock in the lossy JTL demonstrates oscillatory behavior with leading peaks resembling solitary waves if the losses are weak.
@@ I Introduction
We studied previously the shock waves in the lossy Josephson transmission line (JTL) JTL kogan1 ; kogan2 ; kogan3 and kinks and solitons in the lossless (actually, without any shunting at al) JTL kogan2 ; kogan3 . The present work had several aims. First we would like to analyse the interaction between the ”sound” (small amplitude small wave vector harmonic wave) and the shock wave. Second we would like to establish the relation between the shock waves and the kinks. And third, we would like to additionally study the weak waves, and, in particular, to look for the cases when the nonlinear equation, describing such waves in the JTL, can be integrated in terms of elementary functions.
@@ IV.1 The travelling waves
Excluding $V$ from (IV.1) we obtain
$\overline{U}^{2}\frac{d\varphi}{d\tau}=\frac{d}{d\tau}\left(\sin\varphi+\frac{Z_{J}}{R_{J}}\frac{d\varphi}{d\tau}+\frac{C_{J}}{C}\frac{d^{2}\varphi}{d\tau^{2}}\right).$ (26)
Integrating (26) we get
$\frac{d^{2}\varphi}{d\tilde{\tau}^{2}}+\gamma\frac{d\varphi}{d\tilde{\tau}}+\sin\varphi=\overline{U}^{2}\varphi+F,$ (27)
where $\tilde{\tau}\equiv\tau\sqrt{C/C_{J}}=t/\sqrt{L_{J}C_{J}}$, $\gamma\equiv\sqrt{L_{J}/C_{J}}\big{/}R_{J}$ is the damping coefficient, and $F$ is the constant of integration. Equation (27) reminds the equation
@@
Taking into account the boundary conditions (23) and the equation for the shock speed (9), we obtain that $F$ in Eq. (27) is
$F=\frac{\varphi_{1}\sin\varphi_{2}-\varphi_{2}\sin\varphi_{1}}{\varphi_{1}-\varphi_{2}}.$ (29)
Hence Eq. (27) can be written down as
@@
Because of the invariance of the system when all phases are shifted by $2\pi$, it is enough to consider $\varphi_{1}\in(-\pi,\pi)$. Because of the condition (6), we should consider only $\varphi_{1}\in(-\pi/2,\pi/2)$, and because the physics is obviously symmetric with respect to simultaneous inversion of all phases $\varphi\to-\varphi$, in the following we consider explicitly (everywhere apart from Fig. 7) only $\varphi_{1}\in(0,\pi/2)$.
@@
We considered previously the potential maximum at $\varphi_{2}$ case only in the absence of shunting kogan2 . We called such travelling waves the kinks. Now we understand that similar kinks exist also in the lossy JTL (for $-\varphi_{1}<\varphi_{2}<\varphi_{1}/2$). Looking at Fig. 5 (right), presenting the potential for the kink, we realize, that since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $\varphi_{1}$, $\varphi_{2}$ and $\gamma$ should satisfy definite relation. For weak kinks such relation will be obtained in section IV.5. Thus in the absence of losses ($\gamma=0$), only the kinks with $\varphi_{2}=-\varphi_{1}$, are possible kogan2 .
@@
$\overline{U}^{2}-\overline{u}^{2}\left(\varphi_{2}\right)=0.$ (34)
@@
Following the venerable tradition to tell the same story twice, we will now present an approach to integration of the equation (26), alternative to that we used above. Introducing the voltage at the Josephson junction
$E=\frac{d\varphi}{d\tilde{\tau}}$ (35)
as the new dependent variable, considering $\varphi$ as the independent variable and dividing both parts by the common multiplier $E$, we reduce the order of Eq. (26) and write it down as
$\frac{1}{2}\frac{d^{2}E^{2}}{d\varphi^{2}}+\gamma\frac{dE}{d\varphi}=\overline{U}^{2}-\overline{u}^{2}(\varphi).$ (36)
This equation will be used in Section IV.5 to study weak kinks(shocks).
Note that in the framework of the quasi-continuum approximation E is (up to a numerical multiplier) just the voltage on the JJ. Thus Eq. (36) gives the voltage on the JJ as the function of the Josephson phase.
IV.3 The shocks speed vs. the kinks speed
@@
From the inequalities (37) follows that a shock can not split into two shocks moving in the same direction. Actually we can make even stronger statement: two shocks moving in the same direction will merge. In fact, let there is the first shock $\varphi_{0}\leftarrow\varphi_{1}$ and the second shock $\varphi_{2}\leftarrow\varphi_{0}$ behind it. Because of inequalities (37) the speed of the second shock is larger, and the speed of the first shock is smaller than $\overline{u}(\varphi_{0})$. The statement is proved. Note that due to a one-dimensional nature of our problem we don’t have to consider the corrugation instability of the shock wave dyakov1 ; lubchich ; semenko ; robinet ; bates .
@@
IV.4 The quasi-solitons within the shocks
We studied previously the kinks and the solitons in the absence of damping kogan2 . In Section IV.2 we have shown that the kinks exist also in the presence of damping. So what about the solitons?
Looking at the Fig. 4 we understand that in the absence of damping ($R_{J}=\infty\rightarrow\gamma=0$) the motion of the fictitious particle starting and ending at $\varphi=\varphi_{1}$ is possible. This corresponds to the boundary conditions (23) with $\varphi_{2}=\varphi_{1}$ (the soliton). When we switch on the damping the solitons don’t exist any more, because the particle can’t return to its initial position. However, for weak damping, the particle after leaving the initial equilibrium position $\varphi_{1}$ will nearly return to this position after reflection from the opposite wall of the potential well (see Fig. 4). The part of the shock described by that motion will look very much like the soliton. The point of the reflection $\varphi_{0}$ can be found from the equation
$\Pi(\sin\varphi_{0})\approx\Pi(\sin\varphi_{1}).$ (39)
Note that the leading peak will be followed by the other ones, only (because of the continuous decrease of energy of the particle) the distance between the successive peaks will be decreasing and each successive peak will look less and less like the real soliton.
IV.5 Weak kinks and elementary weak shocks
For weak wave, characterized by the condition $\varphi_{1}-\varphi_{2}\ll 1$, the r.h.s. of (30) can be approximated as
@@ IV.5 Weak kinks and elementary weak shocks
and (30) can be simplified to the damped Helmholtz-Duffing (dHD) equation
@@
For $\varphi_{2}<-\varphi_{3}$, Eq. (43) describes the kink, for $-\varphi_{3}<\varphi_{2}$ the shock.
@@
$\varphi(\tilde{\tau})=\overline{\varphi}+\frac{\Delta\varphi}{2}\tanh\left(\beta\tilde{\tau}\right),$ (45)
where $\Delta\varphi=\varphi_{1}-\varphi_{2}$ and
$\beta=\sqrt{\frac{\alpha}{8}}\Delta\varphi=\sqrt{\frac{\cos\overline{\varphi}}{48}}\Delta\varphi.$ (46)
Note that for $\varphi_{2}<-\varphi_{3}$, Eq. (44) is both the condition of the kink existence (fine tuning we talked about previously) and of the kink being given by the elementary function. The shock doesn’t demand find tuning, so for $-\varphi_{3}<\varphi_{2}$, Eq. (44) is only the condition of the shock being given by the elementary function.
If the assumption $\Delta\varphi\ll 1$ is strengthened to $\Delta\varphi\ll\tan\overline{\varphi}$, Eq. (43) can be approximated as
$\varphi_{\tilde{\tau}\tilde{\tau}}+\gamma\varphi_{\tilde{\tau}}=\frac{\sin\overline{\varphi}}{2}(\varphi-\varphi_{1})(\varphi-\varphi_{2}).$ (47)
It is shown in Appendix D the solution of (47) satisfying the boundary conditions (23) is
$\varphi(\tilde{\tau})=\varphi_{2}+\frac{\Delta\varphi}{\left[\exp(\beta^{\prime}\tilde{\tau})+1\right]^{2}},$ (48)
@@
$\beta^{\prime}=\sqrt{\Delta\varphi\sin\overline{\varphi}/12},\hskip 28.45274pt\gamma=5\beta^{\prime}.$ (49)
In the vicinity of $\varphi=0$ and for $\Delta\varphi\ll\tan\overline{\varphi}$, the condition (44) corresponds to strongly overdamped oscillator, and the condition (49), to slightly overdamped oscillator. The shock described by Eq. (47) exists for any value of $\gamma$. Equation (49) is only the condition of it being given by the elementary function.
In both particular cases studied above, because of the overdamped nature of the oscillations, the shock wave is monotonic. Note that the detailed study of the shock wave structure in a strongly nonlinear lattice with viscous dissipation was presented in Ref. nesterenko where, in particular, the dichotomy between the oscillatory and the monotonic shock waves was analysed quantitatively. We, however, postpone a similar analysis for the shocks in the JTL until the next publication.
It is interesting to see how the results obtained above can be recovered in the framework of the approach based on Eq. (36). If $\varphi_{1}$ and $\varphi_{2}$ are close enough to each other, we can approximate $\cos\varphi$ between $\varphi_{1}$ and $\varphi_{2}$ as a second degree polynomial in $\varphi$
$\cos\varphi=\cos\overline{\varphi}-\varphi^{\prime}\sin\overline{\varphi}-\frac{\varphi^{\prime 2}}{2}\cos\overline{\varphi},$ (50)
where $\varphi^{\prime}=\varphi-\overline{\varphi}$. The boundary conditions (23) obviously give
$E\left(\varphi_{1,2}\right)=0,$ (51)
thus we may try the solution of (36) in the form
$E(\varphi)=\psi\left(\varphi-\varphi_{1}\right)\left(\varphi-\varphi_{2}\right),$ (52)
where $\psi$ is a constant. Substituting (52) into (36) we obtain that (52) is indeed the solution, provided Eq. (44) is valid and $\psi^{2}=\cos\overline{\varphi}/12$. To find $\varphi(\tilde{\tau})$ we have to solve equation
$\frac{d\varphi}{d\tilde{\tau}}=\sqrt{\frac{\cos\overline{\varphi}}{12}}\left(\varphi-\varphi_{1}\right)\left(\varphi-\varphi_{2}\right).$ (53)
The solution is Eq. (45).
When $\Delta\varphi\ll\overline{\varphi}$ we can keep in the r.h.s. of (50) only the first two terms and present (36) (with the accepted precision) as
$\frac{1}{2}\frac{d^{2}E^{2}}{d\varphi^{2}}+\gamma\frac{dE}{d\varphi}=\overline{U}^{2}-\cos\overline{\varphi}+\left(\varphi-\varphi_{2}\right)\sin\overline{\varphi}.$ (54)
In this case the solution of (54) satisfying the boundary conditions Eq. (51) is
$E=\chi\left(\varphi-\varphi_{2}\right)\left[\left(\varphi-\varphi_{2}\right)^{1/2}-\left(\Delta\varphi\right)^{1/2}\right].$ (55)
Substituting (55) into (54) and equating coefficients before the same powers of $\left(\varphi-\varphi_{2}\right)^{1/2}$ in the l.h.s and in the r.h.s. of the equation, we obtain (49) and $\chi^{2}=\sin\overline{\varphi}/3$. The function $E(\varphi)$ being found, we can find the function $\varphi(\tilde{\tau})$ by solving equation
$\frac{d\varphi}{d\tilde{\tau}}=\sqrt{\frac{\sin\overline{\varphi}}{3}}\left(\varphi-\varphi_{2}\right)\left[\left(\varphi-\varphi_{2}\right)^{1/2}-\left(\Delta\varphi\right)^{1/2}\right].$ (56)
The solution is given by (48).
@@ V Conclusions
We established the relation between the shocks existing in the lossy JTL and the kinks which, as we now understand, exist both in the lossy and in the lossless JTL. We also have shown that the profile of the shock in the lossy JTL demonstrates oscillatory behavior with leading peaks resembling solitary waves if the losses are weak.
We found that the nonlinear equation describing the weak kinks and the weak shocks can be integrated (in particular cases) in terms of elementary functions.
@@ Appendix A The JTL composed of superconducting grains
$m\frac{d^{2}y_{n}}{d\tau^{2}}=$displaystyle-- $\frac{\partial}{\partial y_{n}}\left[U\left(y_{n-1}-y_{n}\right)+U\left(y_{n+1}-y_{n}\right)\right]$ (59)
-- $\gamma\frac{dy_{n}}{d\tau},$
@@
$\frac{d^{2}\varphi_{n}}{d\tau^{2}}+\gamma\frac{d\varphi_{n}}{d\tau}+\sin\varphi_{n}$displaystyle-- $\frac{1}{a^{2}}\left(\varphi_{n-1}-2\varphi_{n}+\varphi_{n+1}\right)$ (60)
$=$ $I/I_{c},$
@@ Appendix B The continuum and the discrete JTL
Lossless JTL clearly corresponds to $R_{J}=\infty$. In addition, while talking about the lossless system in the present paper, we had in mind Eq. (B) containing only the second term in the l.h.s.. Previously kogan2 , we considered unshunted JTL in the quasi-continuum approximation, that is Eq. (B) with only the first term in the l.h.s.. However, because of the similarity of the terms, the kinks and the solitons obtained in the framework of these two considerations are qualitatively very similar. (The quantitative differences do exist. Thus the equation for the ”soliton” velocity (40) is different from the equation for the soliton velocity (36) from Ref. kogan2 ).
@@
Appendix D The elementary particular solutions of the generalized dHD equation
@@
$x_{\tau\tau}+\gamma x_{\tau}=\alpha x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right),$ (70)
where $n$ is an integer, $\gamma$, $\alpha$, $x_{1},x_{3}$ are positive constants.
@@
$-\alpha(x_{1}-x_{3})e^{-nm\tau}w^{n+1}-\left(m^{2}-m\gamma+\alpha x_{1}x_{3}\right)w.$ (72)
@@
$n^{2}m^{2}w_{zz}+n\left[(n+1)m^{2}-\alpha x_{1}x_{3}\right]\frac{w_{z}}{z}$
$=\alpha w^{2n+1}-\alpha(x_{1}-x_{3})\frac{w^{n+1}}{z}.$ (75)